### v2

# A try for dark energy in quantum field theory: The vacuum energy of neutrino field

Abstract
In this paper, we try to interpret the origin of dark energy in quantum field theory (QFT). The vacuum energy of a field in QFT is a quartic-divergence input, and its physical contribution to dark energy is calculated in the UV-free scheme. There is a mass-energy window that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected. The normal hierarchical spectrum of neutrinos is preferred by naturalness, and the neutrino mass window set by dark energy is about 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
I Introduction
Physics thrives on crisis Weinberg:1988cp , and the ultraviolet (UV) problem is one of the keys in the development of modern physics. Historically, two dark clouds hung over the classical physics at the end of the 19th century, and then the UV-catastrophe in the dynamical theory of heat became a crisis. The energy quanta was introduced by Planck Planck:1901tja and the UV-catastrophe was solved, and quantum theory was born. In quantum field theory (QFT), the problem of UV divergences reappears, that is, the loop corrections are often UV divergent in a transition process. It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output). How do we extract a finite physical result from a UV-divergent physical input? The paradigm approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator and then removed by counterterms, and this traditional route is based on equivalent transformation and the mathematical structure is $\infty$-$\infty$ - ∞. This paradigm approach is successful in dealing with logarithmic divergences in the standard model (SM). For power-law divergences, there are UV problems, i.e.,
(a) The hierarchy problem of the Higgs mass. Loop corrections to the Higgs mass are power-law divergences, with the fine-tuning of the 125 GeV Higgs.
(b) The non-renormalizable Einstein gravity. For power-law divergences of graviton loops, it requires an infinite number of counterterms when one tries to quantize Einstein gravity.
(c) The cosmological constant (dark energy) problem. The vacuum energy in QFT is a quartic divergence, and its contribution (Planck scale) to dark energy exceeds the observation (a density about $(2.3~{}\mathrm{meV})^{4}$) by 120 orders of magnitude.
How do we deal with the above three UV problems in QFT? The free flow of ideas is important when encountering problems. As pointed by Dirac DiracTheEO , the success of renormalization for logarithmic divergences may be similar to the description of one-electron system with further changes required, and the laws of physics are chosen by nature from possible rules that possess mathematical beauty. Feynman Feynman:1986er noted that mathematical self-consistency and legitimacy are important when dealing with UV divergences. Hence, if we go forward, we might have to consider introducing new ideas and new mathematical techniques into QFT to connect UV divergent inputs to the physical reality. For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an explanation to the three UV problems caused by power-law divergences. For UV divergences of loops, assuming the physical contributions being local and contributions from UV regions being insignificant, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections without UV divergences in calculations, which can reproduce the successful renormalization results for logarithmic divergences in SM. This method takes a new route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$, i.e.,
Input{Equivalent transformation,Analytic $\mathrm{Input}\bigg{\{}\!\!\begin{array}[]{c}\textit{Equivalent transformation}\,,\infty-\infty\\ \textit{Analytic continuation}\,,\,\,\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}{}{}{}\end{array}\!\!\bigg{\}}\mathrm{Output}\,,$ (3)
and there is no UV divergence after the physical analytic continuation $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ in the UV-free scheme. To the power-law divergences of loops, the UV-free scheme provides an alternative interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to the quantization of Einstein gravity (graviton loops) was investigated in Ref. Jia:2024lyy . In this paper, another power-law divergence, dark energy, is of our concern. The cosmological constant (dark energy) problem is an open fundamental challenge in physics today Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Copeland:2006wr . Since the form of vacuum energy ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) of a field is very similar to the integral of a loop, we will try it in the UV-free scheme in the following.
II The vacuum energy in QFT
Dark energy has a negative pressure and plays an important role in the evolution of the universe, whose equation of state is $p_{\mathrm{vac}}=\omega\rho_{\mathrm{vac}}$ with $\omega=-1$ adopted. In QFT, the vacuum energy density $\rho_{\mathrm{vac}}$ can be obtained by summing the zero-point energy of all fields. For a field $i$ with a mass $m_{i}$, its contribution is
$\rho_{0}^{i}=(-1)^{2j}g_{i}\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\sqrt{k^{2}+m_{i}^{2}}\,,$ (4)
where $j$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field. This integral is a quartic divergence, and the form of it is very similar to the integral of a loop. Here the UV-free scheme Jia:2023dub ; Jia:2024lyy is adopted to evaluate the physical contribution of Eq. (4), which is a physical analytic continuation of the integral containing UV divergences in QFT. The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}+m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+m_{i}^{2}\!+\xi}$. The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$, i.e., the physical output, is
$\mathcal{T}_{\mathrm{P}}^{i}\!$ $=$ $\!\Big{[}\mathrm{\int}(d\xi)^{3}\frac{\partial^{3}\mathcal{T}_{\mathrm{F}}^{i}(\xi)}{\partial\xi^{3}}\Big{]}_{\xi\to 0}\!\!+\!C$
$=$ $\!\Big{[}(-1)^{2j}g_{i}\!\mathrm{\int}\!(d\xi)^{3}\!\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{3}{16}\frac{1}{\big{(}\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}\big{)}^{5}}\Big{]}_{\xi\to 0}\!\!+\!C$
$=$ $(-1)^{2j}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log(m_{i}^{2})+C\,\,.$
The result with $C$ set by a typical energy scale $\mu_{\Lambda}$ is
$\mathcal{T}_{\mathrm{P}}^{i}=(-1)^{2j}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}\,.$ (6)
In the limit $m_{i}\to 0$, the result is zero for a massless field. [^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation. Here the energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
III The ripple model and neutrino field
Figure 1: The vacuum energy in the ripple model.
Let us consider the contribution of a field to dark energy. The result of Eq. (6) indicates that if dark energy comes from the vacuum energy, then the contributions of heavy fields should be suppressed by some unknown mechanisms. Here a small model is introduced to give a possible description. Supposing that the typical energy scale $\mu_{\Lambda}$ is a watershed energy scale to a field with a mass $m_{i}$, the contribution of vacuum energy to dark energy is active in the case of $m_{i}\lesssim\mu_{\Lambda}$, while the contribution of vacuum energy to dark energy is inert in the case of $m_{i}\gg\mu_{\Lambda}$. This is the ripple model, as depicted in Fig. 1. Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing about 71% of the critical density $3H_{0}^{2}/8\pi G$, the small dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be about $10-100$ meV. Thus, the dark energy density is sensitive to the neutrino field, especially the neutrino mass. [^2] Here is more discussions about possible contributions from neutrinos, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , etc. Considering Eq. (6) and dark energy density, an estimate of a typical neutrino mass $m_{\nu}^{\mathrm{tp}}$ is obtained, with $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$ is adopted). The vacuum energy of a neutrino field with a mass about $m_{\nu}^{\mathrm{tp}}$ can naturally produces the required dark energy density. As a rough prediction, there should be a neutrino with a mass $\sim$ 10 meV.
Figure 2: The normal hierarchical spectra of neutrinos (normal ordering) with dark energy density considered. The solid curves are masses of neutrinos with the mass of $\nu_{1}$ as the input, and the bands with a 5% change in neutrino mass are plotted for a convenient naturalness estimation. The dotted curves are the values of $\mu_{\Lambda}$ required by the dark energy density $(2.3~{}\mathrm{meV})^{4}$ with neutrino masses inputted. The region between the two dashed lines is the neutrino mass window.
Figure 3: The inverted hierarchical spectra of neutrinos (inverted ordering) with dark energy density considered. The solid curves are masses of neutrinos with the mass of $\nu_{3}$ as the input. The dotted curve is the values of $\mu_{\Lambda}$ required by the dark energy density $(2.3~{}\mathrm{meV})^{4}$.
Now we turn to neutrino masses. The masses of three neutrinos are $m_{1}$, $m_{2}$ and $m_{3}$, and the mass differences between them can be extracted by neutrino oscillations. The mass-splitting results from the Particle Data Group ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering). Considering the naturalness with an effective mass $m_{\nu}^{\mathrm{tp}}$, the quasidegenerate spectrum with $m_{1}\simeq m_{2}\simeq m_{3}\gg\sqrt{|\Delta m_{32}^{2}|}$ is not favored. It is possible to give an estimate on the mass spectrum of neutrinos with $m_{\nu}^{\mathrm{tp}}$, and the results obtained are model-dependent. Here we consider a specific implementation of the ripple model by a test factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ describing the active contribution, and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$. For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe). In this case, the observation of dark energy density indicates that neutrino fields play a major role.
The hierarchical spectra of neutrinos for normal ordering and inverted ordering with dark energy density considered are shown in Fig. 2 and Fig. 3, respectively. In Fig. 3, considering the naturalness of $\mu_{\Lambda}$, it can be seen that the inverted hierarchical spectrum is not favored. Let us look at the normal ordering result shown in Fig. 2. There is a region that the neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required for the dark energy density, i.e., a mass-energy window. Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV. This mass window can be tested by the future experiments.
Here we give a brief discussion about the Hubble tension. A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a Hubble tension. For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~{}\mathrm{meV})^{4}$ and $(2.37~{}\mathrm{meV})^{4}$, respectively. If the energy scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig. 2, the Hubble tension can be explained by this slow-running behavior. If the energy scale $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.
IV Conclusion and Discussion
In this paper, the vacuum energy of a field has been investigated, which is another UV divergence of power-law type in QFT. The active vacuum energy of neutrino fields can naturally produce the required dark energy density in the ripple model, and hence a prediction of a neutrino with a typical mass around 10 meV should be present. Considering the naturalness at an energy scale $\mu_{\Lambda}$, the normal hierarchical spectrum of neutrinos is preferred, and there is a mass-energy window for the neutrino mass spectrum and the dark energy density. The neutrino mass window set by dark energy can be examined by the future experiments. In addition, the value of the energy scale $\mu_{\Lambda}$ describing the active contribution of a field here is derived by the observation of dark energy density, and it is taken as an input parameter with naturalness. If this scale is considered from a theoretical point of view, it may be related to the inflationary field, or the origin of the neutrino mass, etc. Moreover, for UV-divergence inputs in QFT, there may be another route besides the routes mentioned in this article. All roads lead to the same goal. With the joint explorations of multiple routes, we hope it will be possible to get close to the harmonious and unified physical structure hidden beneath UV divergences. We look forward to more explorations.
Acknowledgements.
The author would like to thank James M. Cline for helpful discussions. This work was partly supported by the fund of SWUST under the contract no. 23cxtd59.

### v4

@@
 Abstract
-In this paper, we try to interpret the origin of dark energy in quantum field theory (QFT).
-The vacuum energy of a field in QFT is a quartic-divergence input, and its physical contribution to dark energy is calculated in the UV-free scheme.
-There is a mass-energy window that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected.
-The normal hierarchical spectrum of neutrinos is preferred by naturalness, and the neutrino mass window set by dark energy is about 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
+The quartic-divergent vacuum energy poses an ultraviolet (UV) challenge (the cosmological constant problem) in probing the nature of dark energy.
+Here we try to evaluate the contribution of the vacuum energy to dark energy with a method of the UV-free scheme.
+The result indicates that it is not a problem in the UV region but a question of the contributions of heavy fields being suppressed.
+Then, we propose an effective description via scale decoupling.
+The parameter spaces suggest that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected.
+The normal ordering neutrinos are preferred by naturalness, and the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
 I Introduction
 Physics thrives on crisis Weinberg:1988cp , and the ultraviolet (UV) problem is one of the keys in the development of modern physics.
-Historically, two dark clouds hung over the classical physics at the end of the 19th century, and then the UV-catastrophe in the dynamical theory of heat became a crisis.
-The energy quanta was introduced by Planck Planck:1901tja and the UV-catastrophe was solved, and quantum theory was born.
-In quantum field theory (QFT), the problem of UV divergences reappears, that is, the loop corrections are often UV divergent in a transition process.
+In the context of the UV catastrophe in classical physics, the energy quanta were introduced by Planck Planck:1901tja while studying blackbody radiation problems, which led to the birth of quantum theory.
+In quantum field theory (QFT), the problem of UV divergences reappears, that is, loop corrections of a transition process are often UV divergent.
 It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output).
-How do we extract a finite physical result from a UV-divergent physical input?
-The paradigm approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator and then removed by counterterms, and this traditional route is based on equivalent transformation and the mathematical structure is $\infty$-$\infty$ - ∞.
-This paradigm approach is successful in dealing with logarithmic divergences in the standard model (SM).
-For power-law divergences, there are UV problems, i.e.,
+To extract the finite physical result from a UV-divergent input, the traditional approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator based on equivalent transformation and then removed by counterterms, with the mathematical structure of this route being $\infty$-$\infty$ - ∞.
+This traditional approach is successful in dealing with logarithmic divergences in the standard model (SM).
+When we leave the logarithmic region and move on, there are UV problems for power-law divergences, i.e.,
 (a) The hierarchy problem of the Higgs mass. Loop corrections to the Higgs mass are power-law divergences, with the fine-tuning of the 125 GeV Higgs.
-(b) The non-renormalizable Einstein gravity. For power-law divergences of graviton loops, it requires an infinite number of counterterms when one tries to quantize Einstein gravity.
-(c) The cosmological constant (dark energy) problem.
-The vacuum energy in QFT is a quartic divergence, and its contribution (Planck scale) to dark energy exceeds the observation (a density about $(2.3~{}\mathrm{meV})^{4}$) by 120 orders of magnitude.
-How do we deal with the above three UV problems in QFT?
-The free flow of ideas is important when encountering problems.
-As pointed by Dirac DiracTheEO , the success of renormalization for logarithmic divergences may be similar to the description of one-electron system with further changes required, and the laws of physics are chosen by nature from possible rules that possess mathematical beauty.
-Feynman Feynman:1986er noted that mathematical self-consistency and legitimacy are important when dealing with UV divergences.
-Hence, if we go forward, we might have to consider introducing new ideas and new mathematical techniques into QFT to connect UV divergent inputs to the physical reality.
-For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an explanation to the three UV problems caused by power-law divergences.
-For UV divergences of loops, assuming the physical contributions being local and contributions from UV regions being insignificant, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections without UV divergences in calculations, which can reproduce the successful renormalization results for logarithmic divergences in SM.
-This method takes a new route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$, i.e.,
+(b) The non-renormalizable Einstein gravity. When one tries to quantize Einstein gravity, an infinite number of counterterms are required for graviton loops exhibiting power-law divergences.
+(c) The cosmological constant problem.
+The vacuum energy density in QFT is a quartic divergence, and its contribution (Planck scale) exceeds the critical density for a nearly flat universe by 120 orders of magnitude, with the dark energy density $\sim(2.3~{}\mathrm{meV})^{4}$.
+How do we deal with the above three UV problems?
+The traditional approach exhibits UV limitations in handling power-law divergences.
+In this case, the free flow of ideas is important.
+Dirac DiracTheEO viewed renormalization as a mathematical framework requiring refinement to address UV divergences; the physical laws chosen by nature were found to possess mathematical beauty.
+Feynman Feynman:1986er emphasized that mathematical self-consistency and rigor are essential in addressing UV divergences.
+In the vision of Dyson Dyson:1949ha and Schwinger Schwinger , future approaches to loop corrections would be based on finite quantities.
+Wilsonian effective theory Wilson:1971bg decouples scales by encoding UV effects into low-energy parameters via a physical cutoff, enabling systematic organization of loop contributions.
+Hence, if we go forward, we might have to consider introducing new physical ideas and new mathematical techniques to connect UV divergent inputs to the physical reality.
+For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an interpretation to the three UV problems caused by power-law divergences.
+For UV divergences of loops, assuming the physical contributions being local and insensitive to UV regions, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections, which can reproduce the renormalization results for logarithmic divergences in SM.
+This method takes a route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$,
 Input{Equivalent transformation,Analytic $\mathrm{Input}\bigg{\{}\!\!\begin{array}[]{c}\textit{Equivalent transformation}\,,\infty-\infty\\ \textit{Analytic continuation}\,,\,\,\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}{}{}{}\end{array}\!\!\bigg{\}}\mathrm{Output}\,,$ (3)
-and there is no UV divergence after the physical analytic continuation $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ in the UV-free scheme.
-To the power-law divergences of loops, the UV-free scheme provides an alternative interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to the quantization of Einstein gravity (graviton loops) was investigated in Ref.
+and UV divergences are systematically eliminated through $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ without counterterms.
+Loop corrections in this method are constructed from local finite quantities that maintain scale decoupling, while ensuring UV-safe loop contributions through locality-based analytic continuation.
+To power-law divergences of loops, this method provides an interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to Einstein gravity (graviton loops) was investigated in Ref.
 Jia:2024lyy .
-In this paper, another power-law divergence, dark energy, is of our concern.
+Here another power-law divergence, the vacuum energy, is of our concern.
 The cosmological constant (dark energy) problem is an open fundamental challenge in physics today Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Copeland:2006wr .
-Since the form of vacuum energy ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) of a field is very similar to the integral of a loop, we will try it in the UV-free scheme in the following.
+In exploring the nature of dark energy, an unavoidable question is the contribution of the vacuum energy density $\rho_{\mathrm{vac}}$ that corresponds to a cosmological constant ($\Lambda=8\pi G\rho_{\mathrm{vac}}$).
+Since the form of the vacuum energy of a field ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) is similar to the integral of a loop, we will try it in the UV-free scheme in the following.
 II The vacuum energy in QFT
@@
 For a field $i$ with a mass $m_{i}$, its contribution is
-$\rho_{0}^{i}=(-1)^{2j}g_{i}\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\sqrt{k^{2}+m_{i}^{2}}\,,$ (4)
-where $j$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field.
-This integral is a quartic divergence, and the form of it is very similar to the integral of a loop.
+$\rho_{0}^{i}=(-1)^{2j_{i}}g_{i}\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\sqrt{k^{2}+m_{i}^{2}}\,,$ (4)
+where $j_{i}$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field.
+This integral is a quartic divergence.
 Here the UV-free scheme Jia:2023dub ; Jia:2024lyy is adopted to evaluate the physical contribution of Eq.
-(4), which is a physical analytic continuation of the integral containing UV divergences in QFT.
+(4), which is a physical analytic continuation of the integral containing UV divergences.
 The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}+m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+m_{i}^{2}\!+\xi}$.
@@
 $\mathcal{T}_{\mathrm{P}}^{i}\!$ $=$ $\!\Big{[}\mathrm{\int}(d\xi)^{3}\frac{\partial^{3}\mathcal{T}_{\mathrm{F}}^{i}(\xi)}{\partial\xi^{3}}\Big{]}_{\xi\to 0}\!\!+\!C$
-$=$ $\!\Big{[}(-1)^{2j}g_{i}\!\mathrm{\int}\!(d\xi)^{3}\!\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{3}{16}\frac{1}{\big{(}\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}\big{)}^{5}}\Big{]}_{\xi\to 0}\!\!+\!C$
-$=$ $(-1)^{2j}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log(m_{i}^{2})+C\,\,.$
-The result with $C$ set by a typical energy scale $\mu_{\Lambda}$ is
-$\mathcal{T}_{\mathrm{P}}^{i}=(-1)^{2j}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}\,.$ (6)
+$=$ $\!\Big{[}(-1)^{2j_{i}}g_{i}\!\mathrm{\int}\!(d\xi)^{3}\!\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{3}{16}\frac{1}{\big{(}\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}\big{)}^{5}}\Big{]}_{\xi\to 0}\!\!+\!C$
+$=$ $(-1)^{2j_{i}}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log(m_{i}^{2})+C\,\,.$
+The result with $C$ set by a reference energy scale $\mu_{\Lambda}$ is
+$\mathcal{T}_{\mathrm{P}}^{i}=(-1)^{2j_{i}}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}\,.$ (6)
 In the limit $m_{i}\to 0$, the result is zero for a massless field.
 [^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation.
-Here the energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
-III The ripple model and neutrino field
-Figure 1: The vacuum energy in the ripple model.
+The energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
+Here is an interpretation of scales in QFT.
+The physical result of total tree-level and loop-level contributions in a scattering process is independent of the choice of an energy scale $\mu$, and the total physical result can be equivalently described by a tree-level process with an effective coupling at an effective scale $\mu_{\mathrm{eff}}$.
+The energy scale $\mu_{\Lambda}$ of the vacuum energy is similar to the scale $\mu_{\mathrm{eff}}$ in the total physical result, which is different from the free parameter $\mu$ in order-by-order expansions.
+In addition, $\mu_{\mathrm{eff}}$ changes with the scattering energy, while the scale $\mu_{\Lambda}$ describes the reference scale for the ground states of all fields, i.e., all zero-point fields have the same reference scale $\mu_{\Lambda}$.
+Moreover, $\mu_{\Lambda}$ may be a fixed or slow-running scale with the expansion of the universe.
+III The ripple description and neutrino field
+Figure 1: The vacuum energy in the ripple description.
 Let us consider the contribution of a field to dark energy.
 The result of Eq.
-(6) indicates that if dark energy comes from the vacuum energy, then the contributions of heavy fields should be suppressed by some unknown mechanisms.
-Here a small model is introduced to give a possible description.
-Supposing that the typical energy scale $\mu_{\Lambda}$ is a watershed energy scale to a field with a mass $m_{i}$, the contribution of vacuum energy to dark energy is active in the case of $m_{i}\lesssim\mu_{\Lambda}$, while the contribution of vacuum energy to dark energy is inert in the case of $m_{i}\gg\mu_{\Lambda}$.
-This is the ripple model, as depicted in Fig.
+(6) indicates that it is no longer a problem in the UV region.
+The actual question is revealed, that is, the contributions of heavy fields should be suppressed by some unknown mechanisms that are beyond the current understanding of QFT.
+If we analyze the question with Wilsonian framework Wilson:1971bg , to vacuum energy, there may be scale decoupling between microscopic quantity and macroscopic effect.
+We will try to give it an effective description in the view of scale decoupling, that is, the total zero-point energy is the microscopic energy foundation as a background, and quantum fluctuations with cumulative contributions characterized by $\mu_{\Lambda}$ manifest macroscopic effects through fluctuation-carrying fields.
+The scale $\mu_{\Lambda}$ is a watershed scale to a carrier field with a mass $m_{i}$, that is, the contribution of a field’s vacuum energy to dark energy is active, inert in the case of $m_{i}\lesssim\mu_{\Lambda}$, $m_{i}\gg\mu_{\Lambda}$, respectively.
+This picture likes ripples on a lake, necessitating $\mu_{\Lambda}-$characterized fluctuation factor for ground states of fields (e.g., the energy quanta distribution in blackbody radiation research).
+This is the ripple description, as depicted in Fig.
 1.
-Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing about 71% of the critical density $3H_{0}^{2}/8\pi G$, the small dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be about $10-100$ meV.
+Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing 71% of the critical density $3H_{0}^{2}/8\pi G$, the dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be $10-100$ meV.
 Thus, the dark energy density is sensitive to the neutrino field, especially the neutrino mass.
-[^2] Here is more discussions about possible contributions from neutrinos, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , etc.
+[^2] Here is more discussions about possible neutrino contributions, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , the cosmological quantum vacuum energy of axion and neutrino deVega:2023fvw , etc.
 Considering Eq.
-(6) and dark energy density, an estimate of a typical neutrino mass $m_{\nu}^{\mathrm{tp}}$ is obtained, with $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$ is adopted).
-The vacuum energy of a neutrino field with a mass about $m_{\nu}^{\mathrm{tp}}$ can naturally produces the required dark energy density.
-As a rough prediction, there should be a neutrino with a mass $\sim$ 10 meV.
-Figure 2: The normal hierarchical spectra of neutrinos (normal ordering) with dark energy density considered.
+(6) and dark energy density, a typical neutrino mass is $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$).
+As a rough prediction, to produce the required dark energy density, there should be a neutrino with a mass $\sim$ 10 meV.
+Figure 2: The normal ordering neutrinos with the dark energy density $(2.3~{}\mathrm{meV})^{4}$.
 The solid curves are masses of neutrinos with the mass of $\nu_{1}$ as the input, and the bands with a 5% change in neutrino mass are plotted for a convenient naturalness estimation.
-The dotted curves are the values of $\mu_{\Lambda}$ required by the dark energy density $(2.3~{}\mathrm{meV})^{4}$ with neutrino masses inputted.
+The dotted curves are the values of $\mu_{\Lambda}$ required by the dark energy density with neutrino masses inputted.
 The region between the two dashed lines is the neutrino mass window.
-Figure 3: The inverted hierarchical spectra of neutrinos (inverted ordering) with dark energy density considered.
+Figure 3: The inverted ordering neutrinos with the dark energy density $(2.3~{}\mathrm{meV})^{4}$.
 The solid curves are masses of neutrinos with the mass of $\nu_{3}$ as the input.
-The dotted curve is the values of $\mu_{\Lambda}$ required by the dark energy density $(2.3~{}\mathrm{meV})^{4}$.
+The dotted curve is the values of $\mu_{\Lambda}$ required by the dark energy density.
 Now we turn to neutrino masses.
 The masses of three neutrinos are $m_{1}$, $m_{2}$ and $m_{3}$, and the mass differences between them can be extracted by neutrino oscillations.
-The mass-splitting results from the Particle Data Group ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering).
+The mass-splitting results ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering).
 Considering the naturalness with an effective mass $m_{\nu}^{\mathrm{tp}}$, the quasidegenerate spectrum with $m_{1}\simeq m_{2}\simeq m_{3}\gg\sqrt{|\Delta m_{32}^{2}|}$ is not favored.
-It is possible to give an estimate on the mass spectrum of neutrinos with $m_{\nu}^{\mathrm{tp}}$, and the results obtained are model-dependent.
-Here we consider a specific implementation of the ripple model by a test factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ describing the active contribution, and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$.
+It is possible to give an estimate on the mass spectrum of neutrinos.
+Here we consider a specific implementation of the ripple description by a test fluctuation factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ in Gaussian distribution to characterize the active contribution of a field (another possible form $e^{-m_{i}/\mu_{\Lambda}}$ in Boltzmann distribution is not preferred by the naturalness of parameter spaces), and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ ($\rho_{\mathrm{vac}}=\sum_{i}(-1)^{2j_{i}}g_{i}^{\ast}\frac{m_{i}^{4}}{64\pi^{2}}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}$).
 For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe).
-In this case, the observation of dark energy density indicates that neutrino fields play a major role.
-The hierarchical spectra of neutrinos for normal ordering and inverted ordering with dark energy density considered are shown in Fig.
+The hierarchical spectra of neutrinos for normal ordering and inverted ordering with the required dark energy density are shown in Fig.
 2 and Fig.
@@
 In Fig.
-3, considering the naturalness of $\mu_{\Lambda}$, it can be seen that the inverted hierarchical spectrum is not favored.
-Let us look at the normal ordering result shown in Fig.
-2.
-There is a region that the neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required for the dark energy density, i.e., a mass-energy window.
+3, considering the naturalness of $\mu_{\Lambda}$, the inverted ordering is not favored.
+For the normal ordering shown in Fig.
+2, there is a region that neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required by the dark energy density, i.e., the parameter spaces forming a mass-energy window.
 Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV.
 This mass window can be tested by the future experiments.
-Here we give a brief discussion about the Hubble tension.
-A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a Hubble tension.
+Let us give a brief discussion about the Hubble tension.
+A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a possible Hubble tension (the tension may be alleviated Freedman:2024eph ).
 For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~{}\mathrm{meV})^{4}$ and $(2.37~{}\mathrm{meV})^{4}$, respectively.
 If the energy scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig.
-2, the Hubble tension can be explained by this slow-running behavior.
-If the energy scale $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.
+2, the Hubble tension can be explained by this slow-running behavior, i.e., a running cosmological constant.
+If $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.
 IV Conclusion and Discussion
-In this paper, the vacuum energy of a field has been investigated, which is another UV divergence of power-law type in QFT.
-The active vacuum energy of neutrino fields can naturally produce the required dark energy density in the ripple model, and hence a prediction of a neutrino with a typical mass around 10 meV should be present.
-Considering the naturalness at an energy scale $\mu_{\Lambda}$, the normal hierarchical spectrum of neutrinos is preferred, and there is a mass-energy window for the neutrino mass spectrum and the dark energy density.
+In this paper, the vacuum energy’s power-law divergence has been established.
+The physical result in the UV-free scheme indicates that it is not a problem in the UV region.
+Freed from the UV divergence covering, the actual question is that contributions of heavy fields should be suppressed by unknown mechanisms.
+Here we give more interpretations about two common questions as mentioned by Ref.
+Cline:2024klo , i.e., why we choose the value of $\mu_{\Lambda}$ in meV scale, and why we ignore the heavy fields that could contribute.
+To the first question about the value of $\mu_{\Lambda}$, it is currently lacking pathways to derive its specific value from fundamental principles (e.g., the values of particle masses and coupling constants in SM).
+Here the scale $\mu_{\Lambda}$ is taken as an input parameter, and its parameter space is derived by the observation of dark energy density with naturalness.
+This energy scale preserves the spacetime flatness of the present-day universe.
+Theoretically, this scale may link to the inflationary field or neutrino mass origin, etc.
+To the second question, this work explores the $\mu_{\Lambda}-$characterized scale decoupling of vacuum energy from microscopic quantity to macroscopic effect, implying heavy fields remain inert.
+The answer to it is a long exploring journey.
+To account for the dark energy density, a neutrino with a typical mass around 10 meV is expected in the ripple description.
+Considering the naturalness at the scale $\mu_{\Lambda}$, the normal hierarchical spectrum of neutrinos is preferred, and there is a mass-energy window for the neutrino mass spectrum and the dark energy density.
 The neutrino mass window set by dark energy can be examined by the future experiments.
-In addition, the value of the energy scale $\mu_{\Lambda}$ describing the active contribution of a field here is derived by the observation of dark energy density, and it is taken as an input parameter with naturalness.
-If this scale is considered from a theoretical point of view, it may be related to the inflationary field, or the origin of the neutrino mass, etc.
 Moreover, for UV-divergence inputs in QFT, there may be another route besides the routes mentioned in this article.
@@
 Acknowledgements.
-The author would like to thank James M. Cline for helpful discussions. This work was partly supported by the fund of SWUST under the contract no. 23cxtd59.
+The author thanks James M. Cline for useful discussions. This work was partly supported by the fund of SWUST under the contract no. 23cxtd59.

### v5

@@
 The result indicates that it is not a problem in the UV region but a question of the contributions of heavy fields being suppressed.
-Then, we propose an effective description via scale decoupling.
-The parameter spaces suggest that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected.
+Then, we explore an effective description via scale decoupling.
+The parameter spaces suggest that the vacuum energy of active neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV $(10^{-3}$ eV) is expected.
 The normal ordering neutrinos are preferred by naturalness, and the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
@@
 It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output).
-To extract the finite physical result from a UV-divergent input, the traditional approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator based on equivalent transformation and then removed by counterterms, with the mathematical structure of this route being $\infty$-$\infty$ - ∞.
+To extract the finite result from a UV-divergent input, the traditional approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator based on equivalent transformation and then removed by counterterms, with the mathematical structure of this route being $\infty$-$\infty$.
 This traditional approach is successful in dealing with logarithmic divergences in the standard model (SM).
@@
 (a) The hierarchy problem of the Higgs mass. Loop corrections to the Higgs mass are power-law divergences, with the fine-tuning of the 125 GeV Higgs.
-(b) The non-renormalizable Einstein gravity. When one tries to quantize Einstein gravity, an infinite number of counterterms are required for graviton loops exhibiting power-law divergences.
+(b) The non-renormalizable Einstein gravity. When one tries to quantize Einstein gravity, an infinite number of counterterms are required for the graviton loops exhibiting power-law divergences.
 (c) The cosmological constant problem.
@@
 How do we deal with the above three UV problems?
-The traditional approach exhibits UV limitations in handling power-law divergences.
+The traditional approach exhibits limitations in handling power-law divergences, that is, UV divergences under equivalent transformation demand mathematically precarious subtractions via counterterms, especially severe for power-law divergences.
 In this case, the free flow of ideas is important.
-Dirac DiracTheEO viewed renormalization as a mathematical framework requiring refinement to address UV divergences; the physical laws chosen by nature were found to possess mathematical beauty.
+Dirac DiracTheEO regarded renormalization as mathematically provisional in its handling of UV divergences; the physical laws chosen by nature were found to possess mathematical beauty.
 Feynman Feynman:1986er emphasized that mathematical self-consistency and rigor are essential in addressing UV divergences.
 In the vision of Dyson Dyson:1949ha and Schwinger Schwinger , future approaches to loop corrections would be based on finite quantities.
-Wilsonian effective theory Wilson:1971bg decouples scales by encoding UV effects into low-energy parameters via a physical cutoff, enabling systematic organization of loop contributions.
-Hence, if we go forward, we might have to consider introducing new physical ideas and new mathematical techniques to connect UV divergent inputs to the physical reality.
-For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an interpretation to the three UV problems caused by power-law divergences.
-For UV divergences of loops, assuming the physical contributions being local and insensitive to UV regions, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections, which can reproduce the renormalization results for logarithmic divergences in SM.
-This method takes a route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$,
-Input{Equivalent transformation,Analytic $\mathrm{Input}\bigg{\{}\!\!\begin{array}[]{c}\textit{Equivalent transformation}\,,\infty-\infty\\ \textit{Analytic continuation}\,,\,\,\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}{}{}{}\end{array}\!\!\bigg{\}}\mathrm{Output}\,,$ (3)
+Wilsonian effective theory Wilson:1971bg decouples scales via a physical cutoff that encodes UV effects into low-energy parameters.
+Hence, is it not time to explore a route that is intrinsically finite?
+For a new route, it first reproduces the renormalization results for logarithmic divergences, then addresses the three UV problems from power-law divergences.
+For loops, assuming the physical contributions being local and insensitive to UV regions, a method of the UV-free scheme Jia:2023dub is introduced to derive loop contributions, which reproduces the renormalization results for logarithmic divergences.
+It takes a route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$,
+$\mathrm{Input}\bigg\{\!\!\begin{array}[]{c}\textit{Equivalent transformation}\,,\infty-\infty\\ \textit{Analytic continuation}\,,\,\,\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}\end{array}\!\!\bigg\}\mathrm{Output}\,$, (3)
 and UV divergences are systematically eliminated through $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ without counterterms.
-Loop corrections in this method are constructed from local finite quantities that maintain scale decoupling, while ensuring UV-safe loop contributions through locality-based analytic continuation.
-To power-law divergences of loops, this method provides an interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to Einstein gravity (graviton loops) was investigated in Ref.
+The loop contributions constructed through locally finite quantities are intrinsically free of UV divergences, as guaranteed by locality-based analytic continuation.
+For power-law divergences, this method inherently obviates UV subtractions and thus interprets the Higgs mass hierarchy problem within SM Jia:2023dub , and an application to Einstein gravity was investigated in Ref.
 Jia:2024lyy .
-Here another power-law divergence, the vacuum energy, is of our concern.
-The cosmological constant (dark energy) problem is an open fundamental challenge in physics today Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Copeland:2006wr .
-In exploring the nature of dark energy, an unavoidable question is the contribution of the vacuum energy density $\rho_{\mathrm{vac}}$ that corresponds to a cosmological constant ($\Lambda=8\pi G\rho_{\mathrm{vac}}$).
-Since the form of the vacuum energy of a field ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) is similar to the integral of a loop, we will try it in the UV-free scheme in the following.
+Here we focus on the UV-divergent vacuum energy.
+The cosmological constant (dark energy) problem remains a fundamental challenge in physics Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Padmanabhan:2002ji ; Copeland:2006wr ; Li:2011sd .
+In exploring the nature of dark energy, a central and unavoidable question is the vacuum energy density $\rho_{\mathrm{vac}}$ that corresponds to a cosmological constant ($\Lambda=8\pi G\rho_{\mathrm{vac}}$).
+Since the form of the vacuum energy ($\varpropto\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}\!+\!m^{2}}$) resembles a loop integral, the UV-free scheme will be employed as a uniform treatment.
 II The vacuum energy in QFT
@@
 $\rho_{0}^{i}=(-1)^{2j_{i}}g_{i}\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\sqrt{k^{2}+m_{i}^{2}}\,$, (4)
-where $j_{i}$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field.
-This integral is a quartic divergence.
-Here the UV-free scheme Jia:2023dub ; Jia:2024lyy is adopted to evaluate the physical contribution of Eq.
-(4), which is a physical analytic continuation of the integral containing UV divergences.
+where $j_{i}$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom.
+This is a quartic divergence, yet not a loop contribution.
+The UV-free scheme Jia:2023dub ; Jia:2024lyy , a physical analytic continuation of the UV-divergent integral, is adopted to evaluate the contribution of Eq.
+(4).
 The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}\!+\!m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}$.
-The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$, i.e., the physical output, is
+The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$ is
 $\mathcal{T}_{\mathrm{P}}^{i}\!$ $=$ $\!\Big[\mathrm{\int}(d\xi)^{3}\frac{\partial^{3}\mathcal{T}_{\mathrm{F}}^{i}(\xi)}{\partial\xi^{3}}\Big]_{\xi\to 0}\!\!+\!C$
@@
 In the limit $m_{i}\to 0$, the result is zero for a massless field.
-[^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation.
-The energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
+[^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect from bounded vacuum field modes, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation.
+The new energy scale $\mu_{\Lambda}$ is the scale characterizing vacuum energy, analogous to $\Lambda_{\mathrm{QCD}}$ for the strong interaction and $v$ for the electroweak sector.
 Here is an interpretation of scales in QFT.
-The physical result of total tree-level and loop-level contributions in a scattering process is independent of the choice of an energy scale $\mu$, and the total physical result can be equivalently described by a tree-level process with an effective coupling at an effective scale $\mu_{\mathrm{eff}}$.
-The energy scale $\mu_{\Lambda}$ of the vacuum energy is similar to the scale $\mu_{\mathrm{eff}}$ in the total physical result, which is different from the free parameter $\mu$ in order-by-order expansions.
-In addition, $\mu_{\mathrm{eff}}$ changes with the scattering energy, while the scale $\mu_{\Lambda}$ describes the reference scale for the ground states of all fields, i.e., all zero-point fields have the same reference scale $\mu_{\Lambda}$.
-Moreover, $\mu_{\Lambda}$ may be a fixed or slow-running scale with the expansion of the universe.
+In a scattering process, the full physical result (tree and loop levels) is independent of the free parameter $\mu$ used in perturbative expansions; it can be equivalently described by a tree-level process with an effective coupling at the scale $\mu_{\mathrm{eff}}$.
+The vacuum energy scale $\mu_{\Lambda}$ is analogous to this physical scale $\mu_{\mathrm{eff}}$.
+Unlike $\mu_{\mathrm{eff}}$, which varies with scattering energy, $\mu_{\Lambda}$ acts as a common reference scale for all field ground states, and may be fixed or slowly running with cosmic evolution.
 III The ripple description and neutrino field
 Figure 1: The vacuum energy in the ripple description.
-Let us consider the contribution of a field to dark energy.
-The result of Eq.
-(6) indicates that it is no longer a problem in the UV region.
-The actual question is revealed, that is, the contributions of heavy fields should be suppressed by some unknown mechanisms that are beyond the current understanding of QFT.
-If we analyze the question with Wilsonian framework Wilson:1971bg , to vacuum energy, there may be scale decoupling between microscopic quantity and macroscopic effect.
-We will try to give it an effective description in the view of scale decoupling, that is, the total zero-point energy is the microscopic energy foundation as a background, and quantum fluctuations with cumulative contributions characterized by $\mu_{\Lambda}$ manifest macroscopic effects through fluctuation-carrying fields.
-The scale $\mu_{\Lambda}$ is a watershed scale to a carrier field with a mass $m_{i}$, that is, the contribution of a field’s vacuum energy to dark energy is active, inert in the case of $m_{i}\lesssim\mu_{\Lambda}$, $m_{i}\gg\mu_{\Lambda}$, respectively.
-This picture likes ripples on a lake, necessitating $\mu_{\Lambda}-$characterized fluctuation factor for ground states of fields (e.g., the energy quanta distribution in blackbody radiation research).
-This is the ripple description, as depicted in Fig.
-1.
-Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing 71% of the critical density $3H_{0}^{2}/8\pi G$, the dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be $10-100$ meV.
-Thus, the dark energy density is sensitive to the neutrino field, especially the neutrino mass.
-[^2] Here is more discussions about possible neutrino contributions, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , the cosmological quantum vacuum energy of axion and neutrino deVega:2023fvw , etc.
+The QFT vacuum, the ground state of quantum fields, contains zero-point fluctuations from the uncertainty principle.
+Their short-distance energy density is given by Eq.
+(6), showing that a field’s contribution to dark energy is not a UV problem.
+The central question thus becomes how mechanisms beyond current QFT suppress heavy-field contributions.
+The short-distance result cannot be directly identified with the macroscopic cosmological constant, implying a fundamental scale separation.
+In the Wilsonian framework Wilson:1971bg , this signifies a decoupling between microscopic zero-point energy and its macroscopic effect.
+We therefore adopt an effective description based on this scale decoupling, taking the total zero-point energy of free field modes as the physical baseline.
+Physical vacuum energy is then defined as the deviation from this baseline, caused by boundary conditions or interactions.
+We propose that dark energy originates from the cumulative effect of vacuum fluctuations mediated by carrier fields at the characteristic scale $\mu_{\Lambda}$.
+Here $\mu_{\Lambda}$ acts as the filtering scale for a field of mass $m_{i}$: fields with $m_{i}\lesssim\mu_{\Lambda}$ contribute actively, while those with $m_{i}\gg\mu_{\Lambda}$ remain inert.
+This picture is captured by the ripple description (Fig.
+1): just as a net water displacement requires coherent ripples, a net physical contribution requires vacuum fluctuations to achieve macroscopic coherence at the scale set by $\mu_{\Lambda}$ via a fluctuation factor structuring the carrier field, which selects coherent contributions on macroscopic scales, analogous to how the Boltzmann factor selects energy quanta in thermal radiation.
+Given the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy at 71% of the critical density $3H_{0}^{2}/8\pi G$, the dark energy density is about $(2.3~\mathrm{meV})^{4}$, corresponding to $\mu_{\Lambda}\sim 10\!-\!100$ meV, indicating sensitivity to the neutrino field, in particular to neutrino mass.
+[^2] Other phenomenological proposals explore neutrino contributions via, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , the cosmological axion and neutrino deVega:2023fvw , etc.
 Considering Eq.
 (6) and dark energy density, a typical neutrino mass is $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$).
-As a rough prediction, to produce the required dark energy density, there should be a neutrino with a mass $\sim$ 10 meV.
+As a rough prediction, there should be a neutrino with a mass $\sim$ 10 meV.
 Figure 2: The normal ordering neutrinos with the dark energy density $(2.3~\mathrm{meV})^{4}$.
@@
 It is possible to give an estimate on the mass spectrum of neutrinos.
-Here we consider a specific implementation of the ripple description by a test fluctuation factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ in Gaussian distribution to characterize the active contribution of a field (another possible form $e^{-m_{i}/\mu_{\Lambda}}$ in Boltzmann distribution is not preferred by the naturalness of parameter spaces), and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ ($\rho_{\mathrm{vac}}=\sum_{i}(-1)^{2j_{i}}g_{i}^{\ast}\frac{m_{i}^{4}}{64\pi^{2}}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}$).
+Here we consider a specific implementation of the ripple description by a test fluctuation factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ in Gaussian distribution (see the Appendix) to characterize the active contribution of a field, and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ ($\rho_{\mathrm{vac}}=\sum_{i}(-1)^{2j_{i}}g_{i}^{\ast}\frac{m_{i}^{4}}{64\pi^{2}}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}$).
 For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe).
@@
 For the normal ordering shown in Fig.
-2, there is a region that neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required by the dark energy density, i.e., the parameter spaces forming a mass-energy window.
+2, there is a region that neutrino masses at given $\mu_{\Lambda}$ can naturally produce the physical vacuum energy required by the dark energy density, i.e., the parameter spaces forming a mass-energy window.
 Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV.
@@
 For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~\mathrm{meV})^{4}$ and $(2.37~\mathrm{meV})^{4}$, respectively.
-If the energy scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig.
+If the scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig.
 2, the Hubble tension can be explained by this slow-running behavior, i.e., a running cosmological constant.
@@
 IV Conclusion and Discussion
-In this paper, the vacuum energy’s power-law divergence has been established.
-The physical result in the UV-free scheme indicates that it is not a problem in the UV region.
-Freed from the UV divergence covering, the actual question is that contributions of heavy fields should be suppressed by unknown mechanisms.
-Here we give more interpretations about two common questions as mentioned by Ref.
-Cline:2024klo , i.e., why we choose the value of $\mu_{\Lambda}$ in meV scale, and why we ignore the heavy fields that could contribute.
+In this paper, the power-law divergence of vacuum energy is established.
+The physical result obtained in the UV-free scheme indicates it is not a problem in the UV region.
+Freed from the veil of UV divergence, the core question becomes how contributions of heavy fields are suppressed by unknown mechanisms, constituting a new open question.
+This work thereby recasts the cosmological constant problem from a fine-tuning puzzle into a defined physical question, thus clarifying the actual nature of the problem.
+In fundamental physics, such a redefinition is itself an advance, while its full resolution is the next open goal.
+A new framework comes with its necessary transformation of perspective.
+Accordingly, we address two recurrent questions (also noted in Ref.
+Cline:2024klo ): why the value of $\mu_{\Lambda}$ is chosen at the meV scale, and why contributions from heavy fields are neglected.
 To the first question about the value of $\mu_{\Lambda}$, it is currently lacking pathways to derive its specific value from fundamental principles (e.g., the values of particle masses and coupling constants in SM).
-Here the scale $\mu_{\Lambda}$ is taken as an input parameter, and its parameter space is derived by the observation of dark energy density with naturalness.
-This energy scale preserves the spacetime flatness of the present-day universe.
+Here $\mu_{\Lambda}$ is taken as an input parameter, and its parameter space is derived by the observation of dark energy density with naturalness.
 Theoretically, this scale may link to the inflationary field or neutrino mass origin, etc.
+A connection to neutrinos is of particular interest, as their mass scale would interface quantum vacuum physics with the cosmos.
 To the second question, this work explores the $\mu_{\Lambda}$-characterized scale decoupling of vacuum energy from microscopic quantity to macroscopic effect, implying heavy fields remain inert.
-The answer to it is a long exploring journey.
+This suggests that the dark energy problem could be minimally addressed by introducing the new scale $\mu_{\Lambda}$ into SM.
 To account for the dark energy density, a neutrino with a typical mass around 10 meV is expected in the ripple description.
@@
 We look forward to more explorations.
-Acknowledgements.
+Acknowledgments
 The author thanks James M. Cline for useful discussions. This work was partly supported by the fund of SWUST under the contract no. 23cxtd59.
+Appendix A Fluctuation factor
+For a mass $m$ carrier field, the $\mu_{\Lambda}$-characterized fluctuation factor describes macroscopic effects from cumulative vacuum fluctuations.
+Its potential form is derived as follows.
+The first is basic assumptions, i.e.,
+(a) Independence assumption.
+Consider a system with $N$ independent random fluctuations $\{q_{i}\}_{i=1}^{N}$, satisfying: Zero expectation, with $E[q_{i}]=0$; Finite variance, with $Var(q_{i})=\sigma_{i}^{2}$ (bounded fluctuation strength); Statistical independence between random variables $q_{i}$.
+(b) Linear superposition.
+The total cumulative effect $X$ is a linear superposition of individual fluctuations with $X=\sum_{i=1}^{N}a_{i}q_{i}$, where $a_{i}$ being weight coefficients, and the system exhibits no significant nonlinear coupling.
+Next is the derivation via the characteristic function.
+For each $q_{i}$, its characteristic function $\phi_{i}(\tau)$ = $E[e^{i\tau q_{i}}]$ expands to $\phi_{i}(\tau)$ = $1+i\tau E[q_{i}]-\frac{\tau^{2}}{2}E[q_{i}^{2}]+\cdots\approx 1-\frac{\sigma_{i}^{2}\tau^{2}}{2}$, where $\cdots$ denotes higher-order terms.
+By independence, the characteristic function of $X$ is the product of individual characteristic functions, $\phi_{X}(\tau)=\prod_{i=1}^{N}\phi_{i}(a_{i}\tau)\approx\prod_{i=1}^{N}(1-\frac{a_{i}^{2}\sigma_{i}^{2}\tau^{2}}{2})$.
+Taking the logarithm and linearizing (using $\ln(1+x)\approx x$ for $x\to 0$), the result is $\ln\phi_{X}(\tau)\approx-\frac{\tau^{2}}{2}\sum_{i=1}^{N}a_{i}^{2}\sigma_{i}^{2}$ $\Rightarrow$ $\phi_{X}(\tau)=\exp(-\frac{\sigma_{X}^{2}\tau^{2}}{2})$, where $\sigma_{X}^{2}=\sum_{i=1}^{N}a_{i}^{2}\sigma_{i}^{2}$.
+This is the characteristic function of a Gaussian distribution.
+In this paper, we assume that the cumulative effect $X$ of vacuum fluctuations manifests macroscopic effects through carrier fields.
+Specifically, the fluctuation factor $f(m)$ is set by $\phi_{X}(\tau)$, and it can be obtained via inverse Fourier transform of $\phi_{X}(\tau)$ with a matching condition $X\to m$, that is, $f(m)=A\frac{1}{2\pi}\int_{-\infty}^{\infty}e^{-i\tau m}\phi_{X}(\tau)\,d\tau=\frac{A}{\sqrt{2\pi\sigma_{X}^{2}}}\exp(-\frac{m^{2}}{2\sigma_{X}^{2}})=B\exp(-\frac{m^{2}\pi B^{2}}{A^{2}})$, with $B=\frac{A}{\sqrt{2\pi\sigma_{X}^{2}}}$.
+Considering photon field as the carrier field for normalization, one has $f(0)=1$, and thus $B=1$ is obtained.
+Taking $A^{2}/\pi=2\sigma_{X}^{2}=\frac{1}{k}\mu_{\Lambda}^{2}$, the fluctuation factor is written as $f(m)=e^{-km^{2}/\mu_{\Lambda}^{2}}$, and $k\sim 1$ can be adopted for $\mu_{\Lambda}$-characterized fluctuations.
+From another perspective, the macroscopic effect of physical vacuum energy requires a long-range description with characteristic scale $\sim 1/\mu_{\Lambda}$.
+This reflects the filtering of microscopic energy into a macroscopic effective quantity, a principle common to statistical physics.
+A classic example is gas pressure, which stems not from each molecule’s rest energy $mc^{2}$, but from the small fraction in statistically averaged translational kinetic energy.
+Thus, macroscopic observables are only scale-selected subsets of microscopic energies.
+In this analogy, the dark energy density is a macroscopic observable governed by a $\mu_{\Lambda}$-scale filter.
+The microscopic zero-point energy of a quantum field is enormous, yet only fluctuation modes coherent over $\sim 1/\mu_{\Lambda}$ contribute effectively.
+For a field with $m\gg\mu_{\Lambda}$, its coherence length (the Compton wavelength $\lambda_{c}\!\sim\!1/m$) is far smaller than $1/\mu_{\Lambda}$.
+Consequently, a macroscopic volume of size $\sim 1/\mu_{\Lambda}$ contains many independent fluctuation patches with random phases; their contributions cancel under macroscopic averaging via a random-walk process, leading to exponential suppression.
+Thus, the cumulative effect of a heavy field is strongly suppressed.
+The observed small dark energy density therefore does not contradict huge microscopic energies, but shows the cosmos is only sensitive to modes passing the decoupling filter $\lambda_{c}\gtrsim 1/\mu_{\Lambda}$.
+This phase-statistical filtering provides a statistical explanation for the smallness of dark energy density.